# HealthSLM-Bench: Benchmarking Small Language Models for Mobile and Wearable Healthcare Monitoring

### v1

@@ 5.2 Deployment Efficiency
To investigate efficiency and computational cost in real-world deployment, we ran inference with the two top-performing models, Phi-3-mini-4k and TinyLlama-1.1B, which were instructional-tuned using LoRA, on an iPhone 15 Pro Max with 8GB memory capacity. Since the SOTA LLM HealthAlpaca-lora-7b Kim et al. (2023) did not release its checkpoint, we compared the on-device performance of selected SLMs against the baseline Llama-2-7b (the backbone of HealthAlpaca-lora-7b) using PMData to evaluate deployment efficiency. For fair comparison, we random select a total of ten samples from PMData for both Llama-2-7b, Phi-3-mini-4k and TinyLlama-1.1 to evaluate latency and hardware utilization.
@@
Table 6: Efficiency & Utilization of LLMs & SLMs on the HotPotQA and PMData.
@@ Appendix B Small Language Models
  • Phi-3-mini-4k-Instruct Microsoft [2024]: Microsoft’s smallest model in the Phi-3 family. It has 3.8 billion parameters, trained on a combination of synthetic data and selected publicly available website data, with an emphasis on high-quality and reasoning-dense properties.
  • Phi-3.5-mini-Instruct Microsoft [2024]: A upgrade version of phi-3-mini-4k-instruct. It is built in the same architecture and dataset upon phi-3, but trained with a focus on reasoning dense data for better instruction alignment and multi-step reasoning.
  • TinyLlama-1.1B-Chat-v1.0 TinyLlama [2024]: Distilled version of Llama 2. It uses the same architecture and tokenizer as LLaMA but is compact with 1.1 billion parameters. It was fine-tuned on the UltraChat dataset (contains field-cross synthetic dialogues generated by ChatGPT), making it compatible with a wide range of tasks.
  • Gemma2-2B-it Google [2024]: Google’s SOTA open-source model, built on the same research and technology as the Gemini models but scaled down to 2 billion parameters. It is well-suited for text generation tasks such as question answering, summarization, and reasoning.
  • SmolLM-1.7B-Instruct HuggingFaceTB [2024]: HuggingFace’s flagship model, it has 1.7 billion parameters and is trained on SmolLM-Corpus which consists of synthetic textbooks, stories, and educational Python and web samples.
  • Qwen2-1.5B-Instruct Qwen [2024]: Ailibaba’s state-of-the-art SLM in Qwen2 family. It has only 1.5 billion parameters and is trained on diverse instruction-followed tasks. The included coding and mathematics data for training makes it perform well in coding and quantitative reasoning tasks.
  • Qwen2.5-1.5B-Instruct Qwen et al. [2025]: An upgraded version of Qwen2. It is built on the same dataset and architecture, but places greater emphasis on coding and mathematics tasks, making it more optimized for reasoning and math.
  • Llama-3.2-1B-Instruct Meta AI [2024]: Meta-llama’s state-of-the-art SLM. It shares the identical architecture and pre-trained datasets upon Llama3, but is compressed to 1B parameters.
  • Llama-3.2-3B-Instruct Meta AI [2024]: 3B version of Llama-3.2-1B-Instruct.
@@ D.1 Performance Evaluation
Accuracy (%) Bishop [2006] measures the proportion of correctly predicted instances out of all instances. It provides an overview of whether a model performed well overall, with higher values indicating better performance. However, accuracy does not capture the severity or magnitude of errors in misclassified cases, as all errors are treated equally.
Mean Absolute Error (MAE) Hastie et al. [2009] quantifies the average magnitude of prediction errors by computing the absolute difference between predicted and actual values. Lower MAE indicates better alignment with ground truth. Unlike Accuracy, which only reflects correctness, MAE distinguishes between small and large errors. For example, predicting “3” when the true label is “4” yields an error of 1, while predicting “3” when the true label is “10” yields an error of 7. Thus, MAE captures not only whether predictions are correct but also how close incorrect predictions are to the true values.
@@ D.2 Efficiency and Utilization Evaluation
To further evaluate the efficiency and the actual latency in real cases, all state-of-the-art (SOTA) SLMs that show strong promise will be deployed in processing healthcare field data on a real iPhone 15 Pro Max. To better demonstrate the importance of efficiency on mobile devices, the widely used LLM, Llama 2, is selected and serves as a comparison to the fine-tuned SLMs. The following metrics suggested by MobileAIBench Murthy et al. [2023] are adopted to evaluate both efficiency and utilization:

### v2

@@ 5.2 Deployment Efficiency
To investigate efficiency and computational cost in real-world deployment, we ran inference with the two top-performing models, Phi-3-mini-4k and TinyLlama-1.1B, which were instructional-tuned using LoRA, on an iPhone 15 Pro Max with 8GB memory capacity. Since the SOTA LLM HealthAlpaca-lora-7b [23] did not release its checkpoint, we compared the on-device performance of selected SLMs against the baseline Llama-2-7b (the backbone of HealthAlpaca-lora-7b) using PMData to evaluate deployment efficiency. For fair comparison, we random select a total of ten samples from PMData for both Llama-2-7b, Phi-3-mini-4k and TinyLlama-1.1 to evaluate latency and hardware utilization.
@@
Table 6: Efficiency & Utilization of LLMs & SLMs on the PMData dataset.
@@ Appendix B Small Language Models
  • Phi-3-mini-4k-Instruct [32]: Microsoft’s smallest model in the Phi-3 family. It has 3.8 billion parameters, trained on a combination of synthetic data and selected publicly available website data, with an emphasis on high-quality and reasoning-dense properties.
  • Phi-3.5-mini-Instruct [32]: A upgrade version of phi-3-mini-4k-instruct. It is built in the same architecture and dataset upon phi-3, but trained with a focus on reasoning dense data for better instruction alignment and multi-step reasoning.
  • TinyLlama-1.1B-Chat-v1.0 [33]: Distilled version of Llama 2. It uses the same architecture and tokenizer as LLaMA but is compact with 1.1 billion parameters. It was fine-tuned on the UltraChat dataset (contains field-cross synthetic dialogues generated by ChatGPT), making it compatible with a wide range of tasks.
  • Gemma2-2B-it [34]: Google’s SOTA open-source model, built on the same research and technology as the Gemini models but scaled down to 2 billion parameters. It is well-suited for text generation tasks such as question answering, summarization, and reasoning.
  • SmolLM-1.7B-Instruct [35]: HuggingFace’s flagship model, it has 1.7 billion parameters and is trained on SmolLM-Corpus which consists of synthetic textbooks, stories, and educational Python and web samples.
  • Qwen2-1.5B-Instruct [16]: Ailibaba’s state-of-the-art SLM in Qwen2 family. It has only 1.5 billion parameters and is trained on diverse instruction-followed tasks. The included coding and mathematics data for training makes it perform well in coding and quantitative reasoning tasks.
  • Qwen2.5-1.5B-Instruct [59]: An upgraded version of Qwen2. It is built on the same dataset and architecture, but places greater emphasis on coding and mathematics tasks, making it more optimized for reasoning and math.
  • Llama-3.2-1B-Instruct [36]: Meta-llama’s state-of-the-art SLM. It shares the identical architecture and pre-trained datasets upon Llama3, but is compressed to 1B parameters.
  • Llama-3.2-3B-Instruct [36]: 3B version of Llama-3.2-1B-Instruct.
@@ D.1 Performance Evaluation
Accuracy (%) [60] measures the proportion of correctly predicted instances out of all instances. It provides an overview of whether a model performed well overall, with higher values indicating better performance. However, accuracy does not capture the severity or magnitude of errors in misclassified cases, as all errors are treated equally.
Mean Absolute Error (MAE) [61] quantifies the average magnitude of prediction errors by computing the absolute difference between predicted and actual values. Lower MAE indicates better alignment with ground truth. Unlike Accuracy, which only reflects correctness, MAE distinguishes between small and large errors. For example, predicting “3” when the true label is “4” yields an error of 1, while predicting “3” when the true label is “10” yields an error of 7. Thus, MAE captures not only whether predictions are correct but also how close incorrect predictions are to the true values.
@@ D.2 Efficiency and Utilization Evaluation
To further evaluate the efficiency and the actual latency in real cases, all state-of-the-art (SOTA) SLMs that show strong promise will be deployed in processing healthcare field data on a real iPhone 15 Pro Max. To better demonstrate the importance of efficiency on mobile devices, the widely used LLM, Llama 2, is selected and serves as a comparison to the fine-tuned SLMs. The following metrics suggested by MobileAIBench [31] are adopted to evaluate both efficiency and utilization: